OpenStudy (anonymous):

Mona has a triangular flag. She wants to cut out a similar flag that has dimensions one-fourth of the original flag. If the area of the original flag is 320 square inches, what is the area of the new flag that Mona cuts out?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (cwrw238):

if ratio of one dimension = a:b then ratio of areas = a^2 : b^2

OpenStudy (anonymous):

okay. could you explain further

OpenStudy (cwrw238):

well the smaller flag has a dimension 1/4 of large so larger / smaller = 4 / 1 so area larger/smaller = 4^2 / 1^2 or 16/1

OpenStudy (cwrw238):

so proportinate;y 16 320 -- = --- 1 x where x = area of the smaller flag

OpenStudy (anonymous):

20

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

thank you!!! could you help with my next question?

OpenStudy (cwrw238):

ill take a look